# On Iiro Honkala’s contributions to identifying codes

### v2

@@ 5.3 Removing a vertex
Figure 10: A partial representation (for $I=\{2,3,k-1\}$) of the graphs involved in the proof of Theorem 41; codewords are in black.
@@
The proof of this theorem is based on graphs obtained by a collection of paths $P_{r}$ plus two extra vertices, including $v$, linked to the two extremities of each path; Figure 11 illustrates Theorem 44 for $n=17$, $r=3$.
Figure 11: Illustration of Theorem 44 for $r=3$: in both graphs, the black vertices form an optimal $3$-IdC; ${Id_{3}(G)-Id_{3}(G\setminus\{v\})=13-8=5}$.
@@ 5.4 Removing an edge
Note that $k$ can be chosen arbitrarily and is linked to the order $n$ of $G$ and $G\setminus\{e\}$ by the relation $n=(c_{1}r+c_{2})k+r\lceil\log_{2}(k+2)\rceil+(c_{3}r+c_{4})$ where the integer quadruple ($c_{1},c_{2},c_{3},c_{4})$ takes different values in (a1), (a2), (b1) and (b2) above. This means, roughly speaking, that $k$ is a fraction, depending on $r$, of $n$; therefore, Theorem 46 implies that, given $r\geqslant 2$, there is an infinite collection of graphs $G$ of order $n$ and two positive constants $\alpha$ and $\beta$ with $Id_{r}(G)\geqslant\alpha n$ and, after deletion of a suitable edge $e$, $Id_{r}(G\setminus\{e\})\leqslant\beta\log_{2}n$ (or the other way round: $Id_{r}(G)\leqslant\beta\log_{2}n$ and $Id_{r}(G\setminus\{e\})\geqslant\alpha n$). We can see that adding or deleting one edge can lead to quite a drastic difference for the identification numbers.
@@ 7 Variants and related concepts
Acknowledgements. The authors thank the referees for their remarks and suggestions that helped to correct some errors and improve the article.

### v3

@@ 5.3 Removing a vertex
Figure 10: A partial representation (for $I=\{2,3,k-1\}$) of the graphs involved in the proof of Theorem 5.11; codewords are in black.
@@
The proof of this theorem is based on graphs obtained by a collection of paths $P_{r}$ plus two extra vertices, including $v$, linked to the two extremities of each path; Figure 11 illustrates Theorem 5.14 for $n=17$, $r=3$.
Figure 11: Illustration of Theorem 5.14 for $r=3$: in both graphs, the black vertices form an optimal $3$-IdC; ${Id_{3}(G)-Id_{3}(G\setminus\{v\})=13-8=5}$.
@@ 5.4 Removing an edge
Note that $k$ can be chosen arbitrarily and is linked to the order $n$ of $G$ and $G\setminus\{e\}$ by the relation $n=(c_{1}r+c_{2})k+r\lceil\log_{2}(k+2)\rceil+(c_{3}r+c_{4})$ where the integer quadruple ($c_{1},c_{2},c_{3},c_{4})$ takes different values in (a1), (a2), (b1) and (b2) above. This means, roughly speaking, that $k$ is a fraction, depending on $r$, of $n$; therefore, Theorem 5.16 implies that, given $r\geqslant 2$, there is an infinite collection of graphs $G$ of order $n$ and two positive constants $\alpha$ and $\beta$ with $Id_{r}(G)\geqslant\alpha n$ and, after deletion of a suitable edge $e$, $Id_{r}(G\setminus\{e\})\leqslant\beta\log_{2}n$ (or the other way round: $Id_{r}(G)\leqslant\beta\log_{2}n$ and $Id_{r}(G\setminus\{e\})\geqslant\alpha n$). We can see that adding or deleting one edge can lead to quite a drastic difference for the identification numbers.
@@ 7 Variants and related concepts
Acknowledgements. The author Ville Junnila was funded in part by the Academy of Finland grant 338797. The authors thank the referees for their remarks and suggestions that helped to correct some errors and improve the article.